Mathematics 241 – Multivariable Calculus Quiz 1 – September 7, 2007

Your Name:

A) (5) Give parametric equations for the line in R^{3} containing the points (2, 3, −1) and (4, 1, 0).

B) (5) Find a vector in R^{3} that is perpendicular to both of the vectors a = (2, 1, 2) and b = (3, 2, −2).